# SIR: Multi-view Inverse Rendering with Decomposable Shadow for Indoor Scenes

### 25 Feb 2024 v2

@@ 1 Introduction
Figure 1: The general pipeline and components of our method. Given a set of posed multi-view HDR images of an indoor scene, SIR successfully disentangle the scene appearance into 3D neural fields of shape, global and spatially-varying illumination, soft shadows, and SVBRDFs, which can produce convincing results for several applications such as novel view synthesis, free-viewpoint relighting, object insertion, and material replacement.
@@
The second challenge lies in material estimation, specifically in the decomposition of shadow and albedo. Shadows are highly prevalent and noticeable in indoor scenes because of the primary light sources and the intricate geometric occlusions between objects. This necessitates the decomposition of shadows to achieve realistic effects in scene editing applications. Existing multi-view methods for indoor scenes have not yet effectively resolved this problem. NVDIFFREC [12] tends to bake pre-existing shadows into the learned albedo results. Such deprecated shadows in their albedo results can lead to inaccurate rendering during editing. For example, if the scene is illuminated by a light source placed at a different position, the shadows they produce cannot shift accordingly as expected. Another inverse rendering method [8] proposes a single-view inverse rendering method to effectively decompose shadows. However, it relies on supervision with direct shading, excluding shadow effects, which are not available in real-world scenes. A self-supervised shadow decomposition approach [13] is proposed, yet it struggles with the challenge of unknown light source locations. Consequently, accurately estimating shadows without explicit supervision and with unknown light sources remains a significant challenge.
In response to these challenges, we propose SIR (Shadow Inverse Rendering) (see Fig. 1), an innovative inverse rendering method that efficiently decomposes differentiable shadows for indoor scenes without explicit supervision. To address the challenges in representing accurate geometry and detailed lighting, SIR employs an SDF-based neural radiance field for comprehensive scene representation and leverages multi-view HDR images to significantly provide more information for lighting. Then, to overcome the difficulty in decomposing shadows without supervision, we propose a three-stage material estimation process, including albedo initialization with hard shadows, differentiable shadow optimization, and instance-level BRDF regularization. This approach not only elevates SVBRDF quality but also facilitates sophisticated editing capabilities, such as free-view relighting, object insertion, and material replacement. In summary, the main contributions of this work include:
  • We propose SIR, a new inverse rendering framework to effectively handle complex indoor scenes.
  • We present a novel shadow term that, for the first time, decomposes shadow and albedo in multi-view indoor scenes.
@@ 2 Related Works
Figure 2: Overview of our inverse rendering pipeline. The pipeline consists of three phases: 1) In phase 1, we sample a ray with direction $\bm{v}$ and spatial point x from the given posed HDR images. The geometry network $f_{d}$ learns the signed distance $d$, and the HDR-radiance network $f_{c}$ learns radiance $\hat{C}$. Ray marching is then employed to obtain the surface point $\hat{\textbf{x}}$. 2) In phase 2, we sample diffuse incoming light $L_{i,d}$ from environment maps $E$ for learning irradiance $I_{ir}$. We also calculate the specular incoming light $L_{i,s}$ and the pseudo hard shadow $\xi$. 3) In phase 3, hard shadow $S_{hard}$ is learned using $\Theta_{h}$ with pseudo ground truth. We then initialize the parameters of $\Theta_{s}$ using the optimized parameters of $\Theta_{h}$. Instance-level BRDF regularizers are applied, and the whole rendering equation is optimized to update the soft shadow $S_{soft}$, albedo $\hat{A}$, and roughness $\hat{R}$.
@@ 2.1 Neural Scene Representations
Recent progress in neural scene representations has enhanced the recovery of scene geometry and radiance for inverse rendering tasks. Neural radiance field (NeRF) [5] employs a single Multilayer Perceptron (MLP) to process a 5D light field, thereby learning the radiance field for generating new viewpoints. However, the geometric accuracy of NeRF for inverse rendering is constrained due to the inherent ambiguity in volume rendering [14]. To address this, methods like NeuS [15] and VolSDF [16] have been developed, transforming Signed Distance Functions (SDF) [6] into densities for more accurate volume rendering. In addition to these implicit approaches, explicit representation methods [17, 18] have also been proposed to accelerate the rendering process. Among them, DMTet [19] and other DMTet-based solutions [12, 20] utilize a differentiable marching tetrahedral layer to refine surface meshes. Recently, many neural scene representations have been proposed with hybrid methods such as InstantNGP [21] and 3D Gaussian [22], which significantly improve the efficiency and reduce the computational cost. However, these methods often struggle to accurately capture the complex topologies and small-scale structures found in indoor scenes. To effectively reconstruct the geometry of indoor environments, we propose an optimization of topology, lighting, and materials using an SDF-based neural radiance field.
@@ 2.2 Lighting Estimation
In the field of inverse rendering, accurately estimating lighting, particularly in indoor scenes, is a complex and essential task. Most current illumination estimation methods operate on single images, with a primary emphasis on integrating virtual objects into real images rather than making substantial alterations to the scene’s illumination [23, 24, 25]. While traditional methods like a single environment map [26, 27] and spherical lobes [28] have been used, they often neglect spatial variations and high-frequency details in lighting. Recent innovations [29, 9] have attempted to improve 3D lighting representation, but still grapple with challenges like spatial instability and the lack of HDR information. A per-pixel spatially-varying spherical Gaussians (SVSG) lighting representation [30] is proposed to capture high-frequency effects and demonstrate that SGs are superior to spherical harmonics (SH) for depicting lighting details in indoor scenes. Neural-PIL [31] proposes a pre-integrated lighting network based on image-based lighting (IBL), showing better performances on conveying global illumination than SGs and SH. Hence, in our method, we utilize a neural outgoing HDR field to represent the IBL at any spatial point, thereby ensuring a more accurate and detailed depiction of indoor lighting scenarios with a focus on physically accurate HDR lighting prediction.
@@ 2.3 Material Estimation
The rendering equation. The traditional rendering equation [38] is formulated following the principles of energy conservation in physics. It computes the outgoing radiance value $L_{o}$ at surface point $\hat{\textbf{x}}$ with surface normal $\hat{\bm{n}}_{x}$ along view direction $\bm{\omega}_{o}$ by integrating over the hemisphere $\Omega^{+}=\{\bm{\omega}_{i}:\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x}>0\}$, where $\bm{\omega}_{i}$ is the incident light direction:
@@
3 Methods
3.1 Overview
Given a set of posed HDR images of an indoor scene captured under static illumination, our method aims to accurately recover the geometry, global spatially-varying illumination, shadows, and Bidirectional Reflectance Distribution Functions (SVBRDFs), which can be used in editing applications such as free-view relighting and insert objects with accurate shadow. We represent global and spatially varying illumination and geometry as an HDR neural radiance field with SDF, and we utilize MLPs to represent an irradiance field (Sec. 3.2). To handle the ambiguity between shadows and albedo of indoor scenes well, we introduce the hard shadow term based on the HDR neural radiance field, and we propose a shadow field to represent the continuous shadow (Sec. 3.3). To eliminate residual shadow and efficiently improve the quality of SVBRDFs, we present a three-stage material estimation strategy with differentiable shadow and BRDF regularization (Sec. 3.4). The overview of our method is illustrated in Fig. 2.
@@ 3.2 Geometry and Lighting Representaion
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ and $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$ denote the transmittance and alpha value of sampled point, respectively. $\delta_{i}$ is the distance between neighboring sampled points. $P$ is the number of sampled points along a ray. $\sigma_{i}$ and $\textbf{c}_{i}$ are predicted attributes by the HDR-radiance network $f_{c}$. Following VolSDF [16], we transform SDF value d to density $\sigma$ by [16]:
$\sigma(\textbf{x})=\begin{cases}\frac{1}{\beta}\left(1-\frac{1}{2}\exp\left(\frac{f_{d}(\textbf{x})}{\beta}\right)\right),&\text{if }f_{d}(\textbf{x})<0,\\ \frac{1}{2\beta}\exp\left(-\frac{f_{d}(\textbf{x})}{\beta}\right),&\text{if }f_{d}(\textbf{x})\geq 0,\end{cases}$ (4)
where $\beta$ is a learnable parameter to control the sparsity near the surface. Following this parameterization, we can get a high quality smooth surface from the zero level-set of predicted SDF $d_{\Omega}(\textbf{x})$ of the space $\Omega\in\mathbb{R}^{3}$, and surface normal can be computed as the gradient of the predicted SDF w.r.t. surface point $\hat{\textbf{x}}$, i.e., $\hat{\bm{n}}_{x}=\frac{\nabla{d}_{\Omega}(\hat{\textbf{x}})}{\|\nabla{d}_{\Omega}(\hat{\textbf{x}})\|}$.
@@
$L_{i}(\textbf{x})=\int_{\Omega^{+}}E(\textbf{x},\bm{\omega}_{i})(\bm{\omega}_{i}\cdot\bm{n}_{x})\text{d}\bm{\omega}_{i},$ (5)
@@
where $\bm{\omega_{i}}$ is the incident light direction, $\textbf{p}_{x}$ is the intersection position with scene surface along a ray starting from x with a direction $-\bm{\omega}_{i}$ by ray tracing, the incoming lighting at position x is the integrate of radiance along any $\bm{\omega_{i}}$ over hemisphere $\Omega^{+}$. We use a simple convolution operation on an environment map E to calculate diffuse lighting $L_{i,d}$ at any spatial position x, and Monte Carlo Path Tracing to obtain the specular lighting $L_{i,s}$.
We introduce the irradiance field by using an MLP $\Theta_{ir}(\textbf{x})$ to represent the irradiance ${I}_{ir}$ at any input 3D location x, which can efficiently model a compact and continuous spatially varying and global illumination of indoor scenes under complex lighting conditions. The irradiance field $I_{ir}(\textbf{x})=\Theta_{ir}(\textbf{x})$ is supervised by $L_{i,d}(\textbf{x})$, which is derived from the outgoing HDR radiance field. Here, utilizing arbitrary spatial positions as inputs, instead of restricting to object surface points, allows for capturing a continuous and spatially varied lighting environment. This strategy is particularly advantageous for applications that involve inserting objects into a scene, as it ensures accurate global illumination across any point in the space. Therefore, the irradiance MLP can effectively simulate realistic lighting conditions, offering consistent and precise illumination regardless of the object’s placement within the environment. It can also significantly enhance the fidelity of the lighting model in complex spatial applications.
@@ 3.3 Hard Shadow Estimation
Shadow term depicts the occlusion between geometry and environment lighting. The phenomenon of cast shadow is significant in indoor scenes where the primary light source is relatively strong, and we note that gradient optimization tends to bake shadows into the albedo [20]. To enable high-quality relighting, we first consider separating hard shadows from scenes to avoid ambiguity with albedo in material estimation. The traditional graphics method using ray marching can directly calculate the hard shadow term according to the relationship between the main light and geometry if the light location and geometry are known. However, formulating the shadow term is still a challenging problem due to the unknown location of the light source.
According to the HDR radiance attribute, HDR can save more accurate radiance value without cut-off compared LDR, and HDR intensity of the main light is much greater than the environment light. Here, we assume that if the incident light intensity from any direction at a spatial point is lower than an intensity threshold, the point will be considered in a hard shadow area, which means $\xi(\textbf{x})=0$. Therefore, we propose a simple strategy to address the hard shadow estimation under the unknown location of the light source derived from the prior HDR radiance field (Sec. 3.2). Formally:
$\xi(\textbf{x})=\begin{cases}{1},&\text{if }\Gamma_{max}(\textbf{x})\geq\mu,\\ {0},&\text{if }\Gamma_{max}(\textbf{x})<\mu,\end{cases}$ (7)
$\Gamma_{max}(\textbf{x})=\max_{i}{\Gamma_{i}(\textbf{x},\bm{\omega}_{i})},\text{where}\ i=1,2,...,N$ (8)
@@
The hard shadow field $S_{hard}(\textbf{x})=\Theta_{h}(\textbf{x})$ is supervised by $\xi(\textbf{x})$, which caches the shadow distribution for an indoor scene to avoid duplicate computation during material estimation. Meanwhile, the hard shadow field can provide the initial shadow value to ensure the accurate gradient flow of material.
@@ 3.4 BRDF
Inspired by  [44] and [12], we leverage the Monte Carlo rendering algorithm and image-based lighting representation to recover albedo $\hat{A}$, roughness $\hat{R}$, and soft shadow $\hat{S}$. The rendering equation Eq. 1 can be written as:
$L_{o}(\hat{\textbf{x}},\bm{\omega}_{o})$ $=\int_{\Omega^{+}}E(\hat{\textbf{x}},\bm{\omega}_{i})(\frac{\hat{A}_{x}}{\pi}+f_{s}(\hat{\textbf{x}},\bm{\omega}_{i},\bm{\omega}_{o}))(\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x})\text{d}\bm{\omega}_{i}$ (9)
@@
where $\hat{\bm{n}}_{x}$ is the normal at a surface point $\hat{\textbf{x}}$; $\bm{\omega}_{i}$ is light incident direction; $\bm{\omega}_{o}$ is view direction; $f_{s}$ is the specular BRDF term at the surface point based on the Microfacet model. Further, we represent the rendering result with two components for learning, where $L_{o,d}$ is the diffuse component and $L_{o,s}$ is the specular component. The detailed formulation can be found in Sec. A of the Appendix.
Stage 1: Albedo initialization with $S_{hard}$. We adopt the diffuse component based on the Lambertian model [38] to optimize coarse albedo $\hat{A}$ predicted by $f_{r}$. Here, we introduce a hard shadow term at the initial phase of gradient optimization, specifically to prevent the encroachment of hard shadows into the albedo in dimly lit areas. This strategic inclusion is pivotal for maintaining and elevating the accuracy of the albedo, ensuring that the intrinsic coloration of surfaces is faithfully rendered without the distortion of shadow artifacts. Then, we optimize the initial albedo from diffuse component $L_{o,d}(\hat{\textbf{x}})$ by utilizing estimated diffuse lighting $E$ (Sec. 3.2) and hard shadow $\xi$ (Sec. 3.3):
$L_{o,d}(\hat{\textbf{x}})=\frac{\hat{A}_{x}}{\pi}\xi(\hat{\textbf{x}})\int_{\Omega^{+}}E(\hat{\textbf{x}},\bm{\omega}_{i})(\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x})\text{d}\bm{\omega}_{i}$ (11)
@@
Note that we have pre-computed the irradiance and hard shadow terms using respective MLPs $\Theta_{ir}$ and $\Theta_{h}$ to represent, which can effectively minimize the number of unnecessary ray samples in the material estimation process, thereby enhancing computational efficiency, so we can write the diffuse component integral as:
$L_{o,d}(\hat{\textbf{x}})=\frac{\hat{A}_{x}}{\pi}S_{hard}(\hat{\textbf{x}})I_{ir}(\hat{\textbf{x}})$ (12)
where $\hat{\textbf{x}}$ is the surface point, which is the intersection point along the camera ray. The spatially varying illumination and hard shadow are first queried from irradiance field $\Theta_{ir}$ and hard shadow field $\Theta_{s}$.
Stage 2: Differentiable shadow optimization. In stage 1, we incorporated a hard shadow, leading to a coarse albedo approximation. However, this method struggles with capturing detailed shadows, largely because the binary nature of the hard shadow field oversimplified the light-shadow interaction. This simplification, combined with shadow detail depending on ray sampling density, can not effectively represent the nuanced intensity of soft shadows, especially those affected by changes in light source intensity and object placement. We tackle this problem by introducing a differentiable soft shadow $S_{soft}(\textbf{x})$ using an MLP to represent the soft shadow field $\Theta_{s}(\textbf{x})$, which is designed to provide a more nuanced representation of shadows, extending beyond the binary constraints of the hard shadow field. Crucially, the initial parameters of this soft shadow field $\Theta_{s}$ are inherited from the hard shadow field $\Theta_{h}$. We simply replace the hard shadow with the differentiable soft shadow in Eq. 12, and update the predicted soft shadow by optimizing the diffuse radiance:
$L_{o,d}(\hat{\textbf{x}})=\frac{\hat{A}_{x}}{\pi}{S_{soft}}(\hat{\textbf{x}})I_{ir}(\hat{\textbf{x}})$ (13)
Stage 3: Instance-level BRDF regularization. We note that lighting and materials are inherently entangled in rendered images, posing a challenge for their complete separation. This can lead to ambiguities during optimization, where elements of lighting could be inaccurately attributed to the albedo. To alleviate the ambiguity, we propose a novel albedo regularizer $\mathcal{L}_{albedo}$. We assume that the luminance of the albedo is uniformly distributed at the instance level, showing no luminance variations. Formally:

### 09 Apr 2024 v3

@@ Abstract
Keywords:
Inverse Rendering Shadow Estimation Material Estimation Scene Editing
@@ 1 Introduction
Figure 1: Given posed multi-view HDR images of an indoor scene, SIR successfully disentangle the scene appearance into inherent attributes, which can produce convincing results for several free-viewpoint editing applications.
@@
The second challenge lies in material estimation, specifically in the decomposition of shadow and albedo. Shadows are highly prevalent and noticeable in indoor scenes because of the primary light sources and the intricate geometric occlusions between objects. This necessitates the decomposition of shadows to achieve realistic effects in scene editing applications. Existing multi-view methods for indoor scenes have not yet effectively resolved this problem.  [36] tends to bake pre-existing shadows into the learned albedo results. Such deprecated shadows in their albedo results can lead to inaccurate rendering during editing. For example, if the scene is illuminated by a light source placed at a different position, the shadows they produce cannot shift accordingly as expected. On the other hand,  [29] proposes a single-view inverse rendering method to effectively decompose shadows. However, it relies on supervision with direct shading, excluding shadow effects, which are not available in real-world scenes.  [25] solves shadow decomposition through self-supervision but cannot handle unknown light source positions. Consequently, accurately estimating shadows without explicit supervision under unknown light sources remains a significant challenge.
To address these challenges, we propose SIR (Shadow Inverse Rendering) (see Fig. 1), which efficiently decomposes shadows for indoor scenes without explicit supervision. SIR employs an SDF-based neural radiance field for scene representation and leverages multi-view HDR images to provide more information for lighting. Then, for unsupervised shadow decomposition, we propose a three-stage material estimation, including albedo initialization with hard shadows, albedo refinement with differentiable soft shadows, and roughness refinement. We introduce instance-level regularization to refine BRDF attributes (albedo and roughness). This approach not only elevates SVBRDF quality but also facilitates editing capabilities. In summary, the main contributions of this work include:
  • We propose SIR, a new inverse rendering framework with a novel shadow term to decompose shadow and albedo in multi-view indoor scenes effectively.
@@ 2.1 Neural Scene Representations
Recent progress in neural scene representations has enhanced the recovery of scene geometry and radiance for inverse rendering tasks. Neural radiance field (NeRF) [34] employs a single Multilayer Perceptron (MLP) to process a 5D light field, thereby learning the radiance field for generating new viewpoints. However, the geometric accuracy of NeRF for inverse rendering is constrained due to the inherent ambiguity in volume rendering [54]. To address this, methods like NeuS [47] and VolSDF [49] have been developed, transforming Signed Distance Functions (SDF) [37] into densities for more accurate volume rendering. In addition to these implicit approaches, explicit representation methods [31, 3] have also been proposed to accelerate the rendering process. Among them, DMTet [41] and other DMTet-based solutions [36, 15] utilize a differentiable marching tetrahedral layer to refine surface meshes. Recently, many neural scene representations have been proposed with hybrid methods such as InstantNGP [35] and 3D Gaussian splatting [20], which significantly improve the efficiency and reduce the computational cost. However, these methods often struggle to accurately capture the complex topologies and small-scale structures in indoor scenes. To effectively reconstruct the geometry of indoor scenes, we propose an optimization of topology, lighting, and materials using an SDF-based neural radiance field.
@@ 2.2 Lighting Estimation
For inverse rendering, accurately estimating lighting, particularly in indoor scenes, is a complex and essential task. Most current illumination estimation methods operate on single images, with a primary emphasis on integrating virtual objects into real images rather than making substantial alterations to the scene’s illumination [19, 11, 52]. While traditional methods like a single environment map [10, 24] and spherical lobes [12] have been used, they often neglect spatial variations and high-frequency details in lighting. Recent innovations [42, 43] have attempted to improve 3D lighting representation, but still grapple with challenges like spatial instability and the lack of HDR information.  [28] propose per-pixel spatially-varying spherical Gaussians (SVSG) lighting representation to capture high-frequency effects and demonstrate that SGs are superior to spherical harmonics (SH) for depicting lighting details in indoor scenes. Neural-PIL [2] proposes a pre-integrated lighting network based on image-based lighting (IBL), showing better performances on conveying global illumination than SGs and SH. Hence, we utilize a neural HDR-radiance field to represent the IBL at any spatial point, thereby ensuring a more accurate and detailed depiction of indoor lighting scenarios with a focus on physically accurate HDR lighting prediction.
@@ 2.3 Material Estimation
3 Methods
3.1 Overview
Given a set of posed HDR images of an indoor scene, our method aims to accurately recover the geometry, global illumination, shadows, and spatially-varying bidirectional reflectance distribution functions (SVBRDFs) (see Fig. 2), which can be used in editing applications with accurate shadow. We use the SDF to represent global illumination and geometry as an HDR-radiance field, and the additional MLPs to represent an irradiance field (Sec. 3.3). To handle the ambiguity between shadows and albedo well, we introduce a hard shadow term based on the HDR neural radiance field and propose a shadow field to represent the continuous shadows (Sec. 3.4). To eliminate residual shadow and efficiently improve the quality of SVBRDFs, we propose a three-stage material estimation strategy with differentiable shadows and BRDF regularization (Sec. 3.5).
Figure 2: The pipeline consists of three phases: 1) We sample a ray with direction $\bm{v}$ and spatial point x from the given posed HDR images. The geometry network $f_{d}$ learns the signed distance $d$ to obtain surface point $\hat{\textbf{x}}$, and the HDR-radiance network $f_{c}$ learns radiance $\hat{C}$. 2) We obtain the diffuse incoming lighting $L_{i,d}$ by integrating incident radiance from environment maps $E$ for learning irradiance $I$. 3) Hard shadow $S_{hard}$ is learned using $\Theta_{h}$ with pseudo ground truth $\xi$. We then initialize the parameters of $\Theta_{s}$ using the optimized parameters of $\Theta_{h}$. Instance-level BRDF regularizers are applied, and the whole rendering equation is optimized to update $\hat{A}$, $\hat{R}$, and $S_{soft}$.
3.2 Background
The traditional rendering equation [16] is formulated following the principles of energy conservation in physics. It computes the outgoing radiance $L_{o}$ at surface point $\hat{\textbf{x}}$ with surface normal $\hat{\bm{n}}_{x}$ along view direction $\bm{\omega}_{o}$ by integrating over the hemisphere $\Omega^{+}=\{\bm{\omega}_{i}:\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x}>0\}$, where $\bm{\omega}_{i}$ is the incident light direction:
@@ 3.3 Geometry and Lighting Representaion
where $T_{i}=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j})$ and $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$ denote the transmittance and alpha value of sampled point, respectively. $\delta_{i}$ is the distance between neighboring sampled points. $P$ is the number of sampled points along a ray. $\sigma_{i}$ and $\textbf{c}_{i}$ are predicted by the HDR-radiance network $f_{c}$. Following VolSDF [49], we convert the SDF value d to density $\sigma$ and obtain high-quality smooth surface from the zero level-set of predicted SDF in space $\Omega\in\mathbb{R}^{3}$, where the surface normal is computed as the gradient of SDF w.r.t. surface point $\hat{\textbf{x}}$, i.e., $\hat{\bm{n}}_{x}=\frac{\nabla{d}_{\Omega}(\hat{\textbf{x}})}{\|\nabla{d}_{\Omega}(\hat{\textbf{x}})\|}$.
@@
$L_{i,d}(\textbf{x})=\int_{\Omega^{+}}E(\textbf{x},\bm{\omega}_{i})(\bm{\omega}_{i}\cdot\bm{n}_{x})\text{d}\bm{\omega}_{i},$ (4)
@@
where $\bm{\omega_{i}}$ is the incident radiance direction, $\textbf{p}_{x}$ is the intersection position with scene surface along a ray starting from x with a direction $-\bm{\omega}_{i}$ by ray tracing. Here, we use a simple convolution operation on the environment map E to calculate diffuse incoming lighting $L_{i,d}$ at any spatial position x.
To capture a continuous and spatially varied lighting environment, we further introduce an irradiance field $\Theta_{I}(\textbf{x})$ to learn the irradiance ${I}$ at any input 3D location x, supervised by pre-computed incoming lighting $L_{i,d}$. Instead of restricting to surface points, $\Theta_{I}$ utilizes arbitrary spatial positions as inputs, efficiently modeling a compact and continuous spatially varying and global illumination of indoor scenes. It can also significantly enhance the fidelity of the lighting model in complex spatial applications.
@@ 3.4 Hard Shadow Estimation
Shadow depicts the occlusion between geometry and ambient lighting. The phenomenon of cast shadow is significant in indoor scenes with relatively strong primary light sources. We note that gradient optimization during inverse rendering tends to bake shadows into the albedo [15]. To enable high-quality relighting, we first consider decomposing hard shadows from scenes to avoid ambiguities with the albedo during material estimation. The traditional graphics method uses ray marching to calculate hard shadows according to the relative locations between the main light source and geometry. However, modeling shadows is a challenge for inverse rendering where light positions are unknown.
We observe that in HDR environments, the light intensity in areas around light sources and in non-light source areas can differ by several times or even by orders of magnitude. Therefore, we can set $\mu$ as the intensity threshold to distinguish between light and non-light source areas. Here, we assume that if the incident light intensity from any direction at a spatial point is lower than $\mu$, the point will be considered in a hard shadow area, which means $\xi=0$. The incident light intensity can be derived from the pre-trained HDR-radiance field (Sec. 3.3). Formally:
$\xi=\begin{cases}{1},&\text{if }\Gamma_{max}(\textbf{x})\geq\mu,\\ {0},&\text{if }\Gamma_{max}(\textbf{x})<\mu,\end{cases}$ (6)
where $\Gamma_{max}(\textbf{x})=\max_{i}{\Gamma_{i}(\textbf{x},\bm{\omega}_{i})},\ i=1,2,...,N$ and $\Gamma_{i}$ is incoming radiance intensity at spatial position x along $\bm{\omega}_{i}$, which is the sum of radiance (along RGB channels) from the environment map $E(\textbf{x},\bm{\omega}_{i})$. N is the number of sample rays. Since the radiance RGB values of each channel range from 0 to 1 for non-light source areas, we set $\mu=3$ to distinguish light and non-light areas. Thus, this simple strategy can help to simulate hard shadows at unknown light locations. It not only takes into account the complexity of multiple light sources but also recovers shadows caused by intense indirect lighting.
@@ 3.5 BRDF
Inspired by [58] and [36], we leverage the Monte Carlo rendering algorithm and image-based lighting representation to recover albedo $\hat{A}$, roughness $\hat{R}$, and soft shadow $S_{soft}$. The rendering equation (Eq. 1) can be written as:
$L_{o}(\hat{\textbf{x}},\bm{\omega}_{o})$ $=\int_{\Omega^{+}}E(\hat{\textbf{x}},\bm{\omega}_{i})(\frac{\hat{A}}{\pi}+f_{s}(\hat{\textbf{x}},\bm{\omega}_{i},\bm{\omega}_{o}))(\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x})\text{d}\bm{\omega}_{i}$ (7)
@@
where $\hat{\bm{n}}_{x}$ is normal at surface point $\hat{\textbf{x}}$, $\bm{\omega}_{i}$ is light incident direction, $\bm{\omega}_{o}$ is view direction. According to the Lambertian model [16], albedo $\hat{A}$ contributes to the diffuse component. And roughness $\hat{R}$ affects the specular component $f_{s}$ according to the Microfacet model. Therefore, we can represent the predicted rendering result $\hat{L}_{o}$ with a diffuse term $\hat{L}_{o,d}$ and a specular term $\hat{L}_{o,s}$ (see Eq. 8). The specular $\hat{L}_{o,s}$ is calculated by Monte Carlo integration. More details can be found in Sec. 0.A of the Appendix.
We optimize the parameters of the rendering equation in the following stages:
Stage 1: Albedo initialization with $S_{hard}$. With coarse albedo $\hat{A}$ predicted by $f_{r}$, we introduce a hard shadow term $\xi$ to the diffuse term $L_{o,d}$ as shown in Eq. 9. By incorporating hard shadows, we can ensure that albedo takes on the inherent colors of the surfaces without shadow artifacts.
$L_{o,d}(\hat{\textbf{x}})=\frac{\hat{A}}{\pi}\xi\int_{\Omega^{+}}E(\hat{\textbf{x}},\bm{\omega}_{i})(\bm{\omega}_{i}\cdot\hat{\bm{n}}_{x})\text{d}\bm{\omega}_{i}$ (9)
@@
Note that we have pre-computed the hard shadow and irradiance terms using respective MLPs $\Theta_{h}$ and $\Theta_{I}$, which can effectively minimize the number of unnecessary ray samples in the material estimation process, thereby enhancing computational efficiency. Therefore, we can re-write the predicted diffuse radiance as $\hat{L}_{o,d}=\frac{\hat{A}}{\pi}S_{hard}I$. In particular, the coarse albedo will only be optimized when $S_{hard}=1$, which means the point is not in a hard shaow area.
Stage 2: Albedo refinement with differentiable $S_{soft}$. As mentioned in Sec. 3.4, using hard shadows will introduce noises to the rendering results. Therefore, we propose a differentiable soft shadow $S_{soft}$ using an MLP to represent the soft shadow field $\Theta_{s}(\textbf{x})$, which is designed to provide a more nuanced representation of shadows, extending beyond the binary constraints of the hard shadow field. Crucially, the initial parameters of this soft shadow field $\Theta_{s}$ are inherited from the hard shadow field $\Theta_{h}$. We simply replace the hard shadow with the differentiable soft shadow, i.e., $\hat{L}_{o,d}=\frac{\hat{A}}{\pi}S_{soft}I$, and update the predicted soft shadow by optimizing the diffuse radiance. Besides, since albedo and shadows interact with each other, we add an instance-level regularizer $\mathcal{L}_{albedo}$ to the albedo. We assume that the luminance of the albedo is uniformly distributed at the instance level, showing no luminance variations. Formally:
@@
where $\Phi(\cdot)$ is an operator to obtain the V value in HSV color domain by converting RGB-space to HSV-space, $\Phi_{inv}(\cdot)$ is an inverse operator to convert HSV-space to RGB-space [15], $M_{i}$ is the prior instance segmentation mask, $K$ denotes the instance classes of the indoor scene, $p$ denotes the minibatch of 3D points, and $\odot$ denotes an element-wise product.
Stage 3: Roughness refinement. In the former two stages, the roughness $\hat{R}$ is directly predicted by the BRDF MLP $f_{r}$. Considering roughness has a similar instance-level assumption, we introduce a roughness smooth regularizer $\mathcal{L}_{rough}$, which constrains the roughness on the same instance level to be similar.